An element X reacts with O2 to form a compound X2O5. If 0.35g of the element forms 0.55g of the compound, what is the approximate atomic mass of the element?
1
Expert's answer
2017-06-17T04:05:09-0400
Let x be atomic mass of the element X (in grams per mole). Atomic mass of O is known, it is 16. Then molar mass of the compound must be nearly 2x+5×16=2x+80.
The reaction is 2X+2.5O_2→X_2 O_5, so 2 moles of X produces 1 mole of the compound, therefore x can be found from the following proportion:
2x -0.35g
2x+80 -0.55g
2x/0.35=(2x+80)/0.55→0.55(2x)=0.35(2x+80)→1.1x=0.7x+28→0.4x=28→
→x=28/0.4=70
Answer:
Approximate atomic mass of X is 70.